# SMP: Reusable Score-Matching Motion Priors for Physics-Based Character Control

### v2

@@ 8. Results
In the following experiments we compare the performance of SMP with AMP (peng2021amp), a widely-used adversarial imitation learning method. Following peng2021amp, the reward weights for AMP are set to $w^{\mathrm{prior}}=w^{g}=0.5$. In addition to the standard AMP method, we also compare with a variant of AMP that uses a frozen discriminator to examine whether the learned adversarial motion prior can be effectively reused without further training. First, a control policy and discriminator are trained jointly using AMP, then the trained discriminator is reused to train new control policies on the same task without further finetuning using data from the new policy. Furthermore, we also include a simple tabula-rasa learning baseline (w/o Prior), where the policy is trained solely to maximize the task reward ($w^{g}=1$), without any motion priors.
@@
8.1. One Motion Prior for $100+\mathrm{N}$ Styles
SMP serves as a modular motion style reward model that guides policy training without requiring access to the original motion dataset. This enables the application of various adaptation techniques to our diffusion-based reward model to further shape the motion prior. We train a general 100style-conditioned motion diffusion model $f({\mathbf{x}}^{i},c)$ using the entire 20-hour 100STYLE dataset, where $c$ is a style label. Classifier-free guidance (CFG) can then be applied to reshape this general prior into style-specific motion priors:
$f_{\mathrm{zombie}}=f({\mathbf{x}}^{i},\emptyset)+w_{\mathrm{cfg}}\left(f({\mathbf{x}}^{i},c_{\mathrm{zombie}})-f({\mathbf{x}}^{i},\emptyset)\right),$
where the style-conditioned prediction $f({\mathbf{x}}^{i},c_{\mathrm{zombie}})$ is applied to the unconditional prediction $f({\mathbf{x}}^{i},\emptyset)$ according to the guidance weight $w_{\mathrm{cfg}}$. These adapted priors can be used to train policies for a given task using different behavioral styles. The performance statistics for policies trained with various styles are reported in Table 1. All policies are trained using the same underlying SMP model with different style labels. We find that simply setting the CFG scale to 1.0 is generally sufficient to specialize the 100-style prior into distinct style-specific priors, enabling agents to perform tasks with diverse stylistic behaviors, as shown in Figure 1. While AMP can achieve comparable performance, it requires training different style-specific discriminators using style-specific datasets. In contrast, using fixed pretrained discriminators (AMP-Frozen) is ineffective in producing the desired stylistics behaviors. With AMP-Frozen, we observe that the discriminator’s accuracy drops over the course of policy training, which indicates that the policy is exploiting the reward model. This exploitation often leads to unnatural behaviors. Qualitative comparisons of the various methods are available in the supplementary video.
@@ 8.1. One Motion Prior for $100+\mathrm{N}$ Styles
SMP supports crafting novel motion priors by manipulating the pretrained diffusion model’s outputs, enabling the creation of new styles that are not present in the original without requiring additional training of the prior. As shown in Figure 6, two different styles can be composed by blending their style-conditioned predictions from the pretrained diffusion model to produce a new “AeroPlane + HighKnees” prior. The composite prior is constructed via:
@@
where $M_{\mathrm{upper}}$ and $M_{\mathrm{lower}}$ are binary masks applied to the upper- and lower-body features, respectively. The resulting prior $f_{\mathrm{comp}}$ can be used directly as the SMP reward $r^{\mathrm{smp}}$. Moreover, our proposed generative state initialization (GSI) can also adopt the composite prior to generate initialization states of the new style, enabling effective exploration when training policies for the new style. Using SMP and GSI together with $f_{\mathrm{comp}}$, our framework is able to train an agent that follows task commands while spreading its arms and lifting its knees high, all without relying on any reference motion data of the specific style.
Figure 6. A pretrained 100-style motion prior can also be adapted to synthesize motion priors of new styles. For example, we create a novel “AeroPlane + HighKnees” prior by blending two existing styles in the $\bm{\epsilon}$-space. This crafted prior, which is used for both generative state initialization and the motion prior objective $r^{\mathrm{smp}}$, enables the agent to perform the target location task with a new, agile style, all without requiring any reference data.
@@
Figure 7. Comparison of normalized task returns across motion control tasks. SMPs demonstrate better sample efficiency, potentially due to the more consistently informative guidance provided by the stationary SMP reward function.
8.2. One Motion Prior for Multiple Tasks
One pretrained score-matching motion prior can be reused to train multiple policies for different tasks, such as steering, target location and dodgeball, as show in Table 2. The score-matching motion prior is trained on a subset of the LaFAN1 dataset (harvey2020robust), containing unstructured running behaviors. The resulting policies achieve high task returns while producing natural gaits. As shown in Figures 5(b) and 5(a), the character can dynamically select and transition between suitable gaits from the prior, to maintain both task performance and motion naturalness. The Steering policy automatically executes a backward jog when the facing direction opposes the target velocity and transitions to a forward jog once they align. It also learns lateral gaits, such as side-stepping and cross-stepping, without the need for an explicit motion planner. The policy trained with SMP also discovers human-like strategies in the Target Location task. The character automatically slows down as it approaches the target. When the target is far away, the character transitions to a fast running behavior. These nuanced behaviors arise purely from the stationary score-matching motion prior combined with a simple target-distance task objective. No motion planner is required to explicitly select which gait the character should perform. SMP provides the policy with the flexibility to adapt behaviors in the dataset in order to create new skills for different tasks. In the challenging Dodgeball task, agents trained with the locomotion prior spontaneously develop agile jumping and dodging skills, as shown in Figure 5(c). These behaviors were not present in the original dataset, but they closely resemble real human strategies in dodgeball. In contrast, AMP fails on this task, potentially due to the instability of the adversarial objective when the required dodging skill deviates significantly from the reference dataset, which contains only locomotion motions. Policies trained using AMP with a frozen discriminator further fail to produce natural behaviors or achieve strong task performance across all tasks, indicating that adversarial priors cannot be effectively reused, even for identical tasks. In comparison, SMP allows a single pretrained motion prior to be reused across different policies and tasks. Moreover, as shown by the task return curves in Figure 7, a fixed score-matching motion prior also offers higher sample efficiency, consistently providing stable and informative guidance throughout reinforcement learning.
8.3. Human-Object Interaction Priors
SMP can be applied not only to locomotion patterns but also to model interaction priors that jointly capture both character and object motions. Given a human-object interaction (HOI) dataset, we train an unconditional HOI diffusion model $f({\mathbf{x}}^{i}_{\mathrm{char}},{\mathbf{x}}^{i}_{\mathrm{obj}})$, which simultaneously learns the dynamics of the character motion ${\mathbf{x}}_{\mathrm{char}}$ and the object motion ${\mathbf{x}}_{\mathrm{obj}}$. The object motion is represented by its rotation and position, both expressed in the character’s local coordinate frame. The prior reward $r^{\mathrm{prior}}_{t}$ is computed following the same procedure described in Algorithm 1 to evaluate the naturalness of character–object interaction dynamics. We demonstrate that the score-matching interaction prior, when combined with task rewards, effectively guides policies to accomplish complex, multi-stage tasks, where the agent can walk toward a box, pick it up, and carry it to an arbitrary target location, with natural, coordinated, and physically faithful motions.
8.4. Learning to Get Up
Getting up from arbitrary fallen states is an essential skill for both humans and humanoid agents. However, this task is particularly challenging, as many existing methods rely on auxiliary rewards to suppress overly dynamic or erratic motions. Existing approaches, such as AMP, which adaptively update the motion prior during training, can effectively learn natural get-up behaviors. As shown in Table 3, our method demonstrates that even with a stationary motion prior, SMP can successfully train a robust get-up policy capable of recovering from random fallen states. In Figure 5(e), the character is knocked far away yet is still able to roll over, push itself up with its hands, and successfully get back to standing. More qualitative results are best viewed in the supplementary video.
Table 3. Performance of SMP on object carry and getup tasks. SMP can be extended to model human–object interaction priors and train effective interaction policies. It also guides the policy to learn robust getup skills from arbitrary fallen states, an essential recovery capability for everyday tasks.
@@ 8.4. Learning to Get Up
Figure 8. Learning curves for single-clip imitation tasks over three random seeds. To evaluate the robustness of SMP, we also train the diffusion models used for motion priors with different random seeds for each experimental run. Our framework demonstrates consistently good performance across seeds.
8.5. Skill Emergence under Data Scarcity
Reinforcement learning with distribution-matching objectives naturally allows agents to generalize and develop skill that not present in the dataset. Different from the typical application of SDS with image-based priors on large dataset, SMP can learn from as little as three seconds of motions, while still allowing the agent to adapt transitions and behaviors beyond the reference data. In this speed following experiment, the dataset contains only three motion clips that move at different speeds. But the policy then learns to adapt those motions to move at a wide range of target speeds, as shown in Figure 5(d). The character naturally adjusts their gait to match the target speed, walking at slower paces and shifting into jogging and running as the speed increases. Although the reference dataset only contains motions at three discrete speeds, the policies learn to modulate the frequency and stride within each gait, producing continuous variations that preserve the style of the original motions. Moreover, the policies exhibit smooth and expressive transitions between gaits not present in the dataset, including building from walk to jog, snapping down from run to walk, and bursting from walk into a sprint. In Figure 7, the learning curves show that the motion prior also improves sample efficiency compared to the baseline without any prior rewards (w/o Prior).
(a) Spinkick
(b) Backflip
(c) Cartwheel
(d) Crawl
8.6. Benchmark: Single-Clip Imitation
To further evaluate the effectiveness of SMP at imitating behaviors from reference motion clips, we compare it with AMP (peng2021amp), AMP-Frozen, and SMILING (wu2025diffusing) on a series of single-clip imitation tasks. The policies are trained using the prior reward $r^{\mathrm{smp}}_{t}$ only. Comparison with a motion tracking method, DeepMimic (peng2018deepmimic), is included for reference. Imitation performance is assessed using the position tracking error $e_{t}^{\mathrm{POS}}$. Unlike motion-tracking methods such as DeepMimic, which are explicitly designed for precise replication of reference motions, SMP, AMP, and SMILING focus on imitating the general style of the motions. Consequently, these methods do not synchronize the policy with the reference trajectory. To fairly evaluate all methods, dynamic time warping (DTW) is applied for SMP, AMP, and SMILING to temporally align the simulated motion with the reference using position tracking error as the cost function (dtw_sakoe). Additionally, pose termination is disabled for DeepMimic to ensure consistency, as it does not apply to non-synchronized approaches.
Quantitative results are summarized in Table 4, with learning curves provided in Figure 8. While AMP exhibits strong imitation performance, it requires retaining the reference motion data throughout policy training. In comparison, SMP accurately reproduces a diverse range of skills, including highly dynamic ones, and demonstrates imitation accuracy and sample efficiency comparable to AMP, all without relying on reference data during training. Across skills, SMP consistently outperforms SMILING and AMP-Frozen. AMP-Frozen, which attempts to remove AMP’s data dependency by simply substituting a pre-trained discriminator, proves inadequate for motion imitation and results in severely degraded behaviors. DeepMimic benefits from explicit phase synchronization and attains lower tracking errors on certain motions. However, SMP still performs competitively. DeepMimic’s performance degrades on challenging motions such as spinkick and cartwheel, where disabling pose termination leads some runs to converge to suboptimal local solutions.

### v3

@@ 8. Experiments
In the following experiments we compare the performance of SMP with AMP (Peng et al., 2021), a widely-used adversarial imitation learning method. Following Peng et al. (2021), the reward weights for AMP are set to $w^{\mathrm{prior}}=w^{g}=0.5$. In addition to the standard AMP method, we also compare with a variant of AMP that uses a frozen discriminator. This configuration evaluates whether the learned adversarial motion prior can be effectively reused without further training. First, a control policy and discriminator are trained jointly using AMP. Then the trained discriminator is reused to train new control policies on the same task without further finetuning using data from the new policy. Furthermore, we include a simple tabula-rasa learning baseline (w/o Prior), where the policy is trained solely to maximize the task reward ($w^{g}=1$), without any motion priors.
8.1. One Motion Prior for $100+\mathrm{N}$ Styles
SMP serves as a modular motion-style reward model that guides policy training without requiring access to the original motion dataset. This enables the application of various adaptation techniques to the diffusion-based reward model that can further shape the motion prior. To demonstrate this capability, we train a general 100style-conditioned motion diffusion model $f({\mathbf{x}}^{i},c)$ using the entire 20-hour 100STYLE dataset, where $c$ is a style label. Classifier-free guidance (CFG) can then be applied to reshape this general prior into style-specific motion priors (Ho and Salimans, 2022):
$f_{\mathrm{style}}=f({\mathbf{x}}^{i},\emptyset)+w_{\mathrm{cfg}}\left(f({\mathbf{x}}^{i},c_{\mathrm{style}})-f({\mathbf{x}}^{i},\emptyset)\right),$
where the style-conditioned prediction $f({\mathbf{x}}^{i},c_{\mathrm{style}})$ is applied to the unconditional prediction $f({\mathbf{x}}^{i},\emptyset)$ according to the guidance weight $w_{\mathrm{cfg}}$. These adapted priors can be used to train policies for a given task using different behavioral styles. The performance statistics for policies trained with various styles are reported in Table 1. All policies are trained using the same underlying SMP model conditioned on different style labels. We find that simply setting the CFG scale to 1.0 is generally sufficient to specialize the 100-style prior into distinct style-specific priors, enabling agents to perform tasks with diverse stylistic behaviors, as shown in Figure 1. While AMP can produce comparable results, it requires training different style-specific discriminators using curated style-specific datasets. In contrast, using a fixed pretrained discriminator (AMP-Frozen) is ineffective in producing the desired stylistics behaviors. With AMP-Frozen, we observe that the discriminator’s accuracy drops over the course of policy training, which indicates that the policy is exploiting errors in the discriminator. This exploitation often leads to unnatural behaviors that nonetheless elicit a high score from the discriminator. Qualitative comparisons of the various methods are available in the supplementary video.
@@ 8.1. One Motion Prior for $100+\mathrm{N}$ Styles
SMP supports crafting novel motion priors by composing the diffusion model’s outputs, enabling the creation of new styles that are not present in the original dataset without requiring additional training of the prior. As shown in Figure 6, two different styles can be composed by blending their style-conditioned predictions from the diffusion model to produce new styles, such as “AeroPlane + HighKnees”. The composite prior is constructed via:
@@
where $M_{\mathrm{upper}}$ and $M_{\mathrm{lower}}$ are binary masks applied to the upper- and lower-body features, respectively. The resulting prior $f_{\mathrm{comp}}$ can be used directly as the SMP reward $r^{\mathrm{smp}}$. Moreover, our proposed generative state initialization (GSI) can also adopt the composite prior to generate initial states of the new style, enabling effective exploration when training policies for the new style. Using SMP and GSI together with $f_{\mathrm{comp}}$, our framework is able to train an agent that follows task commands while spreading its arms, according to the “AeroPlane” style, and lifting its knees high, according to the “HighKnees” style, all without relying on any reference motion data of the composite style. We further demonstrate the flexibility of SMP by crafting more expressive and playful styles. For instance, we create an “Elated + FlickLegs” prior that combines the arm-swinging motion of the “Elated” style with exaggerated lateral leg flicks during locomotion, resulting in an energetic and humorous movement pattern. We also compose the “GracefulArms” and “Spin” styles to produce a ballroom-dance-like behavior, characterized by elegant arm motions and continuous rotational movements.
8.2. One Motion Prior for Multiple Tasks
To further demonstrate the reusability of SMP, we show that a pretrained score-matching motion prior can be reused to train multiple policies for different tasks, such as steering, target location and dodgeball, as show in Table 2. The score-matching motion prior is trained on a subset of the LaFAN1 dataset (Harvey et al., 2020), containing unstructured running behaviors. The resulting policies achieve high task returns while exhibiting naturalistic gaits. As shown in Figures 5(c) and 5(b), the character can dynamically transition between different gaits according to changing goals. The Steering policy automatically executes a backward jog when the heading direction is oriented in an opposite direction with respect to the target direction, then the policy transitions to a forward jog once the two directions are aligned. The policy also learns lateral gaits, such as side-stepping and cross-stepping, when the heading direction is approximately orthogonal to the target direction.
Policies trained with SMP also exhibit human-like strategies in the Target Location task. When the target is far away, the character executes a fast running behavior. When the target is near, the character automatically slows down to slower walking gaits. These nuanced behaviors arise purely from the score-matching motion prior combined with a simple target-distance task objective. No motion planner is required to explicitly specify which gait the character should perform in different scenarios.
SMP provides the policy with the flexibility to adapt behaviors in the dataset in order to create new skills for different tasks. In the challenging Dodgeball task, agents trained with the locomotion prior spontaneously develop agile jumping and dodging skills, as shown in Figure 5(a). These behaviors were not present in the original dataset, but they closely resemble real human strategies in dodgeball. In contrast, AMP is unable to learn effective policies for this task, potentially due to the instability of the adversarial objective when the required dodging skill deviates significantly from the reference dataset, which contains only locomotion motions. Furthermore, policies trained using AMP with a frozen discriminator are unable to produce natural behaviors or achieve strong task performance across all tasks. This is because freezing and reusing the AMP discriminator to train new policies introduces severe out-of-distribution inputs for the discriminator. These out-of-distribution inputs can lead to incorrect reward values that cause the policy to adopt behaviors that are ill-suited for the task. This suggests that adversarial priors cannot be effectively reused, even for identical tasks. In comparison, SMP allows a single pretrained motion prior to be reused across different policies and tasks. Moreover, as shown by the task return curves in Figure 7, SMP tends to exhibit more sample-efficient training. This efficiency improvement may in part be attributed to the fixed reward model, which provides more stationary and stable guidance throughout the policy training process compared to the nonstationary reward model from AMP.
@@ 8.2. One Motion Prior for Multiple Tasks
Table 3. Performance of SMP on object carry and getup tasks. SMP can be extended to model human–object interaction priors and train effective interaction policies. SMP can also be used to train robust getup policies that can recover from arbitrary fallen states.
@@
8.3. Human-Object Interaction Priors
To demonstrate the effectiveness of our system beyond locomotion tasks, we apply SMP to create interaction priors that jointly capture both character and object motions. Given a human-object interaction (HOI) dataset, we train an unconditional HOI diffusion model $f({\mathbf{x}}^{i}_{\mathrm{char}},{\mathbf{x}}^{i}_{\mathrm{obj}})$, which jointly models the distribution of character motions ${\mathbf{x}}_{\mathrm{char}}$ and object motions ${\mathbf{x}}_{\mathrm{obj}}$. The object’s motion is represented by its rotation and position, both expressed in the character’s local coordinate frame. The prior reward $r^{\mathrm{prior}}_{t}$ is computed following the same procedure described in Algorithm 1 to evaluate the naturalness of character–object interactions. This object-interaction SMP is combined with task rewards to train policies to accomplish complex, multi-stage Object Carry tasks, where the agent walks to a box, picks it up, and then carries it to an arbitrary target location. The learned policy is able to successfully perform this task using natural, and coordinated whole-body manipulation skills.
We also demonstrate that a Stair Traversal policy can be trained using an interaction prior constructed from an unconditional human–scene interaction diffusion model. The position and orientation of the staircase are encoded in the character’s local coordinate frame and jointly modeled by the diffusion model. As shown in Figure 8, the resulting policy enables the character to climb up and down the staircase with natural motions and stable foot contacts.
Figure 7. Comparison of normalized task returns across motion control tasks. SMPs demonstrate better sample efficiency, potentially due to the more stable and consistent RL provided by the stationary SMP reward function.
8.4. Learning to Get Up
Getting up from arbitrary fallen states is an essential skill for both humans and humanoid agents. Naturalistic and physically plausible recovery behaviors are particularly important for safety and feasibility in real-world robotic systems (Tao et al., 2022; He et al., 2025). Producing such life-like get-up motions is challenging, as unconstrained policies often exploit highly dynamic or erratic movements that may succeed functionally in simulation but appear unnatural. As a result, many existing approaches rely on carefully designed auxiliary rewards or regularization terms to suppress these artifacts and encourage safe and smooth recovery behaviors (He et al., 2025). Our results in Table 3 demonstrate that SMP can successfully train a robust get-up policy capable of recovering from random fallen states by simply combining a body-height objective with a motion-prior objective, without requiring complicated hand-crafted regularization terms. In Figure 5(e), after the character is knocked down, it naturally rolls over, pushes itself up with its hands, and returns to a standing posture. Additional qualitative results are provided in the supplementary video.
8.5. Skill Emergence under Data Scarcity
Reinforcement learning with distribution-matching objectives naturally allows agents to generalize and develop skill that not present in the dataset. In contrast to the typical application of SDS with image-based priors on large dataset, SMP can learn from as little as three seconds of motions, while still providing the agent with the flexibility to adapt and develop new behaviors beyond those in the original dataset. In the Target Speed task, the training dataset contains only three motion clips at distinct speeds. Despite the limited reference motion data, the learned policy is able to adapt these motions to follow a continuous range of target speeds, as shown in Figure 5(d). The character naturally adjusts its gait to match the target speed, walking at low target speeds, and transitioning into jogging and running as the speed increases. Although the reference dataset only contains motions at three discrete speeds, the policies learn to modulate the frequency and stride within each gait, producing continuous variations that preserve the style of the original motions while traveling at a broad range of speeds. Moreover, the policies exhibit smooth and naturalistic transitions between gaits that are not present in the dataset, including building up speed from a walk to a jog, decelerating from a run back to a walk, and bursting from a walk into a sprint.
As shown in Figure 7, the learning curves indicate that incorporating the motion prior significantly improves sample efficiency compared to the baseline trained without any prior reward (w/o Prior). This improvement arises because the score-matching motion prior provides a dense and informative shaping signal that consistently biases exploration toward physically plausible and natural motions. By constraining the policy to remain within high-likelihood regions of the motion manifold during training, SMP reduces the need for trial-and-error exploration of unnatural behaviors, allowing the policy to focus its learning capacity on task-relevant control. As a result, task rewards and motion priors can reinforce each other, leading to faster convergence and more stable learning.
Figure 8. Example of a stair traversal policy trained with an interaction prior. SMP enables the character to traverse multiple stairs with natural motions.
8.6. Benchmark: Single-Clip Imitation
To further evaluate the effectiveness of SMP at imitating behaviors from reference motion clips, we compare our method with AMP (Peng et al., 2021), AMP-Frozen, and SMILING (Wu et al., 2025a) on a series of single-clip imitation tasks. All policies are trained without the task reward $r^{g}$. In addition to comparisons to prior distribution matching methods, we also include comparisons with a motion tracking method, DeepMimic (DM) (Peng et al., 2018a). Imitation performance is assessed using the position tracking error $e_{t}^{\mathrm{POS}}$, as defined in Appendix C. Unlike motion-tracking methods such as DeepMimic, which are explicitly designed for precise replication of reference motions, SMP, AMP, and SMILING focus on imitating the general style of the motions. Therefore, these methods do not directly synchronize the policy with the reference trajectory. To ensure a more fair comparison across all methods, we follow the evaluation procedure of Peng et al. (2021) and apply dynamic time warping (DTW) to SMP, AMP, and SMILING to temporally align the motions from the simulated character before calculating the position tracking error (Sakoe and Chiba, 1978). Furthermore, pose-error termination is disabled for DeepMimic to ensure a consistent evaluation setup, as this early termination strategy is not applicable to non-tracking methods.
Quantitative results are summarized in Table 4, with learning curves provided in Figure 10. While AMP exhibits strong imitation performance, it requires retaining the reference motion data throughout policy training. In comparison, SMP accurately reproduces a diverse range of skills, including highly dynamic behaviors, such as backflip, and exhibits comparable imitation accuracy and sample efficiency to AMP, all without requiring access to the reference data during training. Across skills, SMP consistently outperforms SMILING and AMP-Frozen. AMP-Frozen, which attempts to remove AMP’s data dependency by simply using a fixed pre-trained discriminator, is unable to accurately reproduce the desired motions. The tracking-based DeepMimic benefits from explicit synchronization between the reference motion and the policy, and thereby attains lower tracking errors on certain motions. However, SMP still performs competitively to this tracking-based method. DeepMimic’s performance degrades on some challenging motions such as spinkick and cartwheel, where disabling pose-error termination leads to some training runs converging to suboptimal behaviors.
(a) Spinkick
(b) Backflip
(c) Cartwheel
(d) Crawl
9. Real-World Robotic Applications
In addition to experiments in simulated environments, SMP can be used to train controllers that enable real-world humanoid robots to perform natural, life-like behaviors. Figure 11 shows behaviors from a walking controller is trained entirely in simulation using SMP and then directly deployed on a G1 humanoid robot. Unlike the simulation experiments, the real-world controller has several modified in order to operate effectively in the real world. First, some state information available in simulation is difficult to reliably estimate on the physical robot. Therefore, the policy observes only a partial representation of the state consisting of proprioceptive information that can be recorded through onboard sensors:
$\mathbf{o}_{t}\triangleq\bigl[\bm{\omega}_{t}^{\text{root}}\!,\,{\mathbf{q}}_{t}^{\text{root}}\!,\,{\mathbf{q}}_{t}\!,\,\dot{{\mathbf{q}}}_{t}\bigr]\in\mathbb{R}^{67},$
which includes the root rotation ${\mathbf{q}}_{t}^{\text{root}}\in\mathbb{R}^{6}$, root angular velocity $\bm{\omega}_{t}^{\text{root}}\in\mathbb{R}^{3}$}, local joint rotations ${\mathbf{q}}_{t}\in\mathbb{R}^{29}$, and local joint velocity $\dot{{\mathbf{q}}}_{t}\in\mathbb{R}^{29}$.
@@ 9. Real-World Robotic Applications
Figure 10. Learning curves of different methods on the single-clip imitation tasks. Three models initialized with different random seeds are trained for each method. To evaluate the robustness of SMP, different diffusion models trained with different random seeds are used to train each SMP policy. SMP demonstrates consistent and effective performance across seeds.
(a) Steering
(b) Recovery Behaviors
(c) Spinkick
Figure 11. SMP policies deployed on a Unitree G1 humanoid robot. The learned motor controllers can reproduce natural locomotion gaits, robust recovery behaviors, and agile motion skills in the real world.
To ensure accurate reward and value estimation during training in simulation, we adopt an asymmetric actor-critic framework in which the SMP diffusion model and value function receive additional privileged information that are not provided to the policy (Pinto et al., 2018). The motion features used by the diffusion model to compute the SMP reward $r^{smp}$ are identical to those described in Section 6.1. The critic (i.e., value function) observes additional privileged state information, such as the root height, root linear velocity, and end-effector positions relative to the root. All features are recorded in the robot’s local coordinate frame. To improve the robustness of the policy to variations in the system dynamics, domain randomization is applied during training in simulation (Peng et al., 2018b), which randomized physical parameters, such as mass, friction and restitution. External perturbations are also applied during training to further enhance the policy’s robustness.
In real-world experiments, the robot demonstrates robust locomotion under external perturbations, automatically adjusting its stride to maintain balance and continue moving forward. These recovery behaviors emerge automatically without relying on a runtime motion planner or manually designed heuristics. Additional qualitative results are provided in the supplementary video.
10. Ablation and Sensitivity Analysis
Our system is one of the first frameworks to construct a reusable motion prior based on score-matching, that can achieve high fidelity results comparable to state-of-the-art adversarial imitation learning methods. In this section, we identify the key design decisions that enable more stable training and higher-quality results.
Ensemble Score-Matching
We compare the proposed ensemble score-matching approach (“Random”), which averages multiple SDS evaluations over a fixed set of diffusion timesteps, against the typical strategy used in prior work, which evaluates SDS at a single randomly sampled timestep (“Random”) (Poole et al., 2022; Luo et al., 2024; Wu et al., 2025a). This comparison is conducted on the single-clip imitation tasks using reference motions of varying difficulties, including Run, Cartwheel, and Backflip. As shown in Table 5, the single-timestep SDS objective performs comparably on simple skills, but struggles on more challenging behaviors. For instance, when imitating a backflip, policies trained with a single SDS evaluation tend to collapse to standing still or falling backwards instead of executing a complete flip. In contrast, ensembling SDS evaluations across a fixed set of diffusion timesteps provides a more consistent and informative reward signal, which enables effective policy training even for challenging acrobatic motions.